# Truth with a Twist: The Rhetoric of Persuasion in Professional vs. Community-Authored Fact-Checks

### v1

@@ 1. Introduction and Background
In comparison to professional fact-checking efforts, CNs offer broader coverage–both in the volume of misinformation addressed and in the diversity of topics requiring regional, cultural, or domain-specific expertise (Augenstein et al., 2025). Beyond these dissemination effects, users tend to express greater trust in corrective interventions when they are accompanied by CNs that explicitly articulate why a post is misleading, in contrast to opaque “false information” labels issued by experts (Drolsbach et al., 2024). Additionally, by engaging a substantially larger and more demographically diverse contributor base than traditional fact-checking organizations (Bouchaud and Ramaciotti, 2025; Augenstein et al., 2025), CNs promote a more democratic and transparent model of content governance. This inclusivity exposes readers to a wider range of perspectives and sources, particularly on controversial or geographically salient topics, reinforcing CNs’ potential as a scalable, community-centered approach to social-media accountability (Bouchaud and Ramaciotti, 2025; Augenstein et al., 2025).
@@
Despite these strengths, there is no consensus among researchers on whether crowdsourced fact-checks are more susceptible to biases than professional ones (Augenstein et al., 2025). Some researchers argue that the bridging algorithm—which requires cross-ideological agreement for a note publication—reduces one-sidedness of CNs relative to professional fact-checks (Kangur et al., 2024). Others emphasize the system’s vulnerabilities to coordination and manipulation. For example, in the context of U.S. politics, posts by Republicans have been shown to receive more notes, both overall and published, than posts by Democrats (Renault et al., 2025). Moreover, Draws et al. demonstrate that crowd workers’ assessments can diverge from expert-assigned ground-truth labels depending on their political orientations and broader cognitive or ideological dispositions (Draws et al., 2022). They further find that crowd workers are more likely than experts to regard misleading claims as truthful (Draws et al., 2022). Within this broader debate, some scholars argue that professional fact-checks are generally more objective in style, whereas CNs may employ persuasive or manipulative linguistic strategies, such as appeals to authority or emotionally charged claims (Kankham and Hou, 2025; Augenstein et al., 2025).
@@
Notwithstanding these concerns, no study to date has carried out a systematic comparison between professional and crowdsourced debunks with respect to their use of persuasive language. This gap motivates the present work, which aims to quantify the extent to which Community Notes adhere to their core values [^1] https://communitynotes.x.com/guide/en/contributing/values of avoiding personal opinions, insults, and manipulative behavior. Accordingly, we pose the following research questions: (a) Is the CNs’ self-governing nature able to identify and penalize the use of persuasion techniques in community notes? (b) Do community-written debunks exhibit a higher degree of persuasive language than professionally-authored ones?
@@ 2. Datasets and Methodology
RQ1. Is the CNs’ self-governing nature able to identify and penalize the use of persuasion techniques in community notes? To address this question, we examined two aspects. First, using the complete set of note ratings from the CN website, we assessed whether the degree of note persuasiveness is correlated with the degree of note helpfulness. We operationalised the degree of note persuasiveness as the number of persuasion techniques identified in a given note. To estimate the helpfulness degree, we considered individual user ratings that mark the note as being Not Helpful, Somewhat Helpful, or Helpful. We then mapped each note rating to a numeric value: Not Helpful-¿1, Somewhat Helpful-¿2, or Helpful-¿3. For notes with multiple ratings, we calculated the average rating value. The association between the degree of persuasiveness and helpfulness level is measured using the Spearman correlation coefficient (Wissler, 1905). For each persuasion technique, we additionally compared the distribution of helpfulness scores between notes that use the technique and those that do not, using the Mann–Whitney U test (Test, 2000). To assess the direction and magnitude of the effect, we computed Cliff’s delta ($\delta$), a non-parametric effect size metric (Macbeth et al., 2011).
@@ 3. Results
Table 2 summarizes the results of the one-sided z-tests for per-technique presence comparison between the datasets. Positive z-values indicate higher prevalence in CNs, and p-values below 0.05 indicate statistically significant differences. The results indicate that the prevalence of persuasion techniques in CNs (underlined values) differs markedly depending on the comparison dataset. Compared to DBKF, only a few techniques (Conversation Killer, Repetition, False Dilemma-No Choice, and Whataboutism) are significantly more prevalent in CNs. For most other techniques, prevalence is lower in CNs, contrary to our initial hypothesis. Compared to EUvsDisinfo, much more techniques (Appeal to Hypocrisy, Appeal to Popularity, Appeal to Values, Consequential Oversimplification, Conversation Killer, Doubt, Obfuscation-Vagueness-Confusion, Questioning the Reputation, Red Herring, Straw Man, and Whataboutism) are significantly more prevalent in CNs. This can be attributed to the stricter format and the domain specificity of the EUvsDisinfo debunks.
@@ 4. Discussion and Conclusions
It should be noted, however, that our per-technique comparisons reveal a more complex picture. While CNs are less persuasive than the DBKF fact-checks along most dimensions, they show higher prevalence for several techniques when compared to EUvsDisinfo. This highlights the importance of domain and organisational context: EUvsDisinfo debunks are produced within a narrow geopolitical niche with very stringent editorial oversight, whereas CNs address a far broader thematic and cultural space. The differences therefore underline that rhetorical variation is shaped not only by authorship (crowd vs. professional) but also by institutional aims, editorial norms, and the type of misinformation being countered.
The analysis of rater behaviour further extends existing knowledge of CNs’ self-regulating mechanisms. The overall degree of persuasiveness in a note exhibits only a negligible relationship with its average helpfulness degree, and—contrary to our hypothesis—the association is slightly positive. However, examining the presence of individual techniques reveals that crowd raters are particularly sensitive to penalise certain types of persuasive rhetoric, especially patriotically charged claims (Flag-Waving); statements that exhibit, biases against individuals through association with specific groups (ethnic, racial, political, ideological, etc.; Guilt by Association); and the notes that purposefully misrepresent the original post by addressing an argument that was not made by the post author (Straw Man). The latter crowd reaction aligns with the broader prevalence of NNN flagging in Community Notes, in which users identify notes as unnecessary or irrelevant to the post (Razuvayevskaya et al., 2025). In contrast, the use of other techniques appears to persuade the raters and increase their trust in CNs. This suggests that CN contributors are sensitive to specific problematic rhetorical strategies even if they do not penalise their cumulative presence.

### v2

@@ 1. Introduction and Background
In comparison to professional fact-checking efforts, CNs offer broader coverage–both in the volume of misinformation addressed and in the diversity of topics requiring regional, cultural, or domain-specific expertise (Augenstein et al., 2025). Beyond these scalability advantages, users tend to express greater trust in corrective interventions when they are accompanied by CNs that explicitly articulate why a post is misleading, in contrast to opaque “false information” labels issued by experts (Drolsbach et al., 2024). Additionally, by engaging a substantially larger and more demographically diverse contributor base than traditional fact-checking organizations (Bouchaud and Ramaciotti, 2025; Augenstein et al., 2025), CNs promote a more democratic and transparent model of content governance. This inclusivity exposes readers to a wider range of perspectives and sources, particularly on controversial or geographically salient topics, reinforcing CNs’ potential as a scalable, community-centered approach to social-media accountability (Bouchaud and Ramaciotti, 2025; Augenstein et al., 2025).
@@
Despite these strengths, there is no consensus among researchers on whether crowdsourced fact-checks are more susceptible to biases than professional ones (Augenstein et al., 2025). Some researchers argue that the bridging algorithm—which requires cross-ideological agreement for a note publication—reduces one-sidedness of CNs relative to professional fact-checks (Kangur et al., 2024). Others emphasise the system’s vulnerabilities to coordination and manipulation. For example, in the context of U.S. politics, posts by Republicans have been shown to receive more notes, both overall and published, than posts by Democrats (Renault et al., 2025). Moreover, Draws et al. demonstrate that crowd workers’ assessments can diverge from expert-assigned ground-truth labels depending on their political orientations and broader cognitive or ideological dispositions (Draws et al., 2022). They further find that crowd workers are more likely than experts to regard misleading claims as truthful (Draws et al., 2022). Within this broader debate, some scholars argue that professional fact-checks are generally more objective in style, whereas CNs may employ persuasive or manipulative linguistic strategies, such as appeals to authority or emotionally charged claims (Kankham and Hou, 2025; Augenstein et al., 2025).
@@
Notwithstanding these concerns, no study to date has carried out a systematic comparison between professional and crowdsourced debunks with respect to their use of persuasive language. This gap motivates the present work, which aims to quantify the extent to which Community Notes adhere to their core values [^1] https://communitynotes.x.com/guide/en/contributing/values of avoiding personal opinions, insults, and manipulative behaviour. Accordingly, we pose the following research questions: (a) Is the CNs’ self-governing nature able to identify and penalise the use of persuasion techniques in community notes? (b) Do community-written debunks exhibit a higher degree of persuasive language than professionally-authored ones?
@@ 2. Datasets and Methodology
RQ1. Is the CNs’ self-governing nature able to identify and penalise the use of persuasion techniques in community notes? To address this question, we examined two aspects. First, using the complete set of note ratings from the CN website, we assessed whether the degree of note persuasiveness is correlated with the degree of note helpfulness. We operationalised the degree of note persuasiveness as the number of persuasion techniques identified in a given note. To estimate the helpfulness degree, we considered individual user ratings that mark the note as being Not Helpful, Somewhat Helpful, or Helpful. We then mapped each note rating to a numeric value: Not Helpful-¿1, Somewhat Helpful-¿2, or Helpful-¿3. For notes with multiple ratings, we calculated the average rating value. The association between the degree of persuasiveness and helpfulness level is measured using the Spearman correlation coefficient (Wissler, 1905). For each persuasion technique, we additionally compared the distribution of helpfulness scores between notes that use the technique and those that do not, using the Mann–Whitney U test (Test, 2000). To assess the direction and magnitude of the effect, we computed Cliff’s delta ($\delta$), a non-parametric effect size metric (Macbeth et al., 2011).
@@ 3. Results
Table 2 summarises the results of the one-sided z-tests for per-technique presence comparison between the datasets. Positive z-values indicate higher prevalence in CNs, and p-values below 0.05 indicate statistically significant differences. The results indicate that the prevalence of persuasion techniques in CNs (underlined values) differs markedly depending on the comparison dataset. Compared to DBKF, only a few techniques (Conversation Killer, Repetition, False Dilemma-No Choice, and Whataboutism) are significantly more prevalent in CNs. For most other techniques, prevalence is lower in CNs, contrary to our initial hypothesis. Compared to EUvsDisinfo, much more techniques (Appeal to Hypocrisy, Appeal to Popularity, Appeal to Values, Consequential Oversimplification, Conversation Killer, Doubt, Obfuscation-Vagueness-Confusion, Questioning the Reputation, Red Herring, Straw Man, and Whataboutism) are significantly more prevalent in CNs. This can be attributed to the stricter format and the domain specificity of the EUvsDisinfo debunks.
@@ 4. Discussion and Conclusions
It should be noted, however, that our per-technique comparisons reveal a more complex picture. While CNs are less persuasive than the DBKF fact-checks along most dimensions, they show higher prevalence for several techniques when compared to EUvsDisinfo. This highlights the importance of domain and organisational context: EUvsDisinfo debunks are produced within a niche geopolitical domain with stringent editorial oversight, whereas DBKF addresses a far broader thematic and organisational space. The differences therefore underline that rhetorical variation is shaped not only by authorship (crowd vs. professional) but also by institutional aims, editorial norms, and the type of misinformation being countered.
The analysis of rater behaviour further extends existing knowledge of CNs’ self-regulating mechanisms. The overall degree of note persuasiveness exhibits only a negligible relationship with its average helpfulness degree, and—contrary to our hypothesis—the association is slightly positive. However, examining the presence of individual techniques reveals that crowd raters are particularly sensitive and penalise certain types of persuasive rhetoric, especially patriotically charged claims (Flag-Waving); statements that exhibit, biases against individuals through association with specific groups (ethnic, racial, political, ideological, etc.; Guilt by Association); and the notes that purposefully misrepresent the original post by addressing an argument that was not made by the post author (Straw Man). The latter crowd reaction could represent a particular case of the widely observed NNN (Note Not Needed) flagging phenomenon in CNs, whereby users label notes as irrelevant to the post (Razuvayevskaya et al., 2025). In contrast, the use of other techniques appears to persuade the raters and increase their trust in CNs. This suggests that CN contributors are sensitive to specific problematic rhetorical strategies even if they do not penalise their cumulative presence.